-- AndreaRossi - 18 Oct 2019

INAF queue observing, 2019 Oct 18- Oct 19

Observer: A. Rossi + M. Faccini and A. Giunta from Rome
Telescope Operator: S. Allanson
Support Astronomer: B. Rothberg
Instrument: LUCI

At the sunset is clear. LUCI2 is not mounted

Summary

Night duration (12 degree el twilight) 1:40 - 12:30 UT = 10 h 50m

Observing time: 10h 20min (94%)

Weather time loss: h min (0%)

technical time loss h 40min (6%)

Data Summary

Project_RUN ObjectsSorted ascending Instrument Configuration ExpTime /Time_tot Images Notes
DDT Epifani 1/I Borisov LBC UB/VRI 0.8h   DDT Epifani
Di Criescenzo AndVII LUCI1 J imaging 0.07h   for AndVII J1 completed
Di Criescenzo AndVII LUCI1 Ks imaging 0.3h   for AndVII Ks1 completed
Calibrations Bias LBC       DDT Epifani
Calibrations darks LUCI1       per Giannini IRAS, FUOri, V912
Calibrations darks LUCI1       per Gargiulo
Telluric Rossi Hip 20379 LUCI1 zJ 200 1"   luci1.0194-0198 DDT Rossi
Telluric HIP115428 LUCI1     luci1.0131-0134 for Feruglio. NOTE: first attempt acquisition saturated
Telluric HIP93727 LUCI1 zJ 200 1" 1.17   luci1.0036-0039 per Nardini
Nardini J1712+57 LUCI1 zJ 200 1" 1.17 0.27h luci1.0048-0055 completed
Feruglio S82X _2058 LUCI1 H/G210 1"slit 3h luci1.0140-0184 seeing limited seeing <1"
Standard Epifani Sa95-275 LBC UB/VRI 0.2h   DDT Epifani
Standard Epifani Sa95-275 LBC UB/VRI 0.2h   DDT Epifani
Calibrations Sky Flats LBC-R V_Bessel (PA=0)     DDT Epifani
DDT Rossi SN2018ibb LUCI1 zJ 200 1" 1.0h luci1.0264-0282 DDT Rossi

Details

00:00 Calibrations
  • darks for Giannini IRAS 60.0sx1 LIR,norm luci1.0003-0007
  • darks for Giannini FUOri 5.0sx6 LIR,int luci1.0008-0012
  • darks for Giannini V912 30.0sx2 LIR,int luci1.0013-0017
  • darks for Giannini telluric 12.0s x5 LIR,int luci1.0018-0022
  • darks for Gargiulo 300.0s x5 MER norm luci1.0023-0027
  • darks for Gargiulo telluric 3.0s x5 LIR int luci1.0028-0032

01:23 Slew to telluric HIP93727 for Nardini J1712+57
  • acq luci1.0033-0035 seeing 1". Note used HeI +J filters! changed script and OT for future use
  • sci luci1.0036-0039

01:35 Slew to Nardini J1712+57
  • acq luci1.0040-0047 take three more truslit after adjusting offset
  • seeing 1.1" on guider at 01:54 0.9" on acquisition images
  • sci luci1.0048-0055
  • seeing 0.9" at 02:12 on DIMM, 0.8" on guider

02:14 Slew to Di Criscienzo AndVII J1 and Ks1
  • seeing 1" at 02:20 on DIMM
  • sci luci1.0056-0065 AND 1
  • seeing 1" at 02:30 on DIMM
  • 02:27 sci luci1.0066-0075 AND 2
  • 02:37 sci luci1.0076-0085 AND 3
  • seeing 1" on camera, 0.8" on images at 02:38
  • 02:46 sci luci1.0086-0095 AND 4
  • seeing 1.2" on DIMM, 0.8" on guiding at 02:48
  • seeing 0.6" on images, 0.8" on guiding at 02:52
  • 02:57 sci luci1.0096-0105 AND 5
  • seeing 0.6" on images, 0.8" on guiding at 03:03
  • 03:07 sci luci1.0106-0115 AND 6 image 108 with some elongation due to guiding problem
  • seeing 0.8" on guiding at 03:09

03:19 Seeing 1" on DIMM Slew to Feruglio telluric HIP115428
  • this telluric is too bright in acquisition with OH+z saturates! change script and DO AGAIN
  • acq luci1.0125-0130
  • sci luci1.0131-0134

03:58 Seeing 0.9" on DIMM Slew to Feruglio S82X _2058
  • acq luci1.0135-0139 took other tru slit after adjusting allignement
  • seeing 0.8" on DIMM at 04:10
  • sci luci1.0140-0184
  • seeing 0.6" on DIMM, 0.6" on guider at 04:25
  • RTD Aladin is slow and subtraction gets stuck half of the times. I often need to stop and start Aladin
  • seeing 0.7" on DIMM, 0.6" on guider at 04:45
  • 05:20 execute a second time
  • seeing 1" on guider at 05:20 - 1.1" on DIMM
  • seeing 0.7" on guider at 05:40 - 0.6" on DIMM
  • seeing 0.6" on guider at 06:03 - 0.5" on DIMM
  • 6:20 LUCI GUI stopped abruptly
  • 6:24 start again but another error has occurred during the execution of image 168
  • 6:30 start again
  • seeing 1.0" on DIMM, 0.7" on guider at 06:38
  • seeing 1.0" on DIMM, 0.9" on guider at 07:24

07:42 Slew to Telluric of Rossi (Hip 20379)
  • seeing 0.9" on DIMM, 1" on guider at 07:51
  • acq luci1.0185-0187
  • sci luci1.0188-0191

07:55 Slew to DDT Rossi SN2018ibb TOO EARLY take telluric again!
  • acq luci1.0192-0193
  • sci luci1.0194-0198

08:15 Slew to DDT Rossi SN2018ibb
  • error in script BUUUU wrong readout
  • acq luci1.0259-0263
  • sci luci1.0264-0272
  • not sure object is in field, added offset y -25 arcsec. Lately I checked and everything is fine.
  • seeing 1.1" on DIMM, 1.3" on guider at 08:49
  • 9:21 start again science (same script)
  • sci luci1.0273-0282
  • seeing 1.3" on guider at 09:38
  • sometimes RTD doesn't work (can not subract images)
  • seeing 0.8" on DIMM at 09:56

9:58 Change of instrument, switch to LBC

10:15 Slew to STD SA95-275 of DDT Epifani (2I/Borisov)
  • 10:20 technical problem with the ADsec
  • 10:49 technical problem resolved
  • 10:50 to STD SA95-275
  • 10:51 collimation and copointing
  • 11:03 re do collimation and copointing
  • 11:15 start acquisition on SA95-275
  • seeing 1.2" on DIMM at 11:15
  • 11:22 slew to STD SA100-267
  • 11:25 start acquisition on SA100-267

11:31 slew to 2I/Borisov (DDT Epifani)
  • 11:32 collimation
  • 11:36 start acquisition on target (U-VRI sequence 1)
  • 11:45 start acquisition on target (U-VRI sequence 2)
  • 11:53 start acquisition on target (U-VRI sequence 3)
  • seeing 1.0" on DIMM at 11:55
  • 12:01 start acquisition on target (B-VRI sequence 1)
  • 12:07 start acquisition on target (B-VRI sequence 2)
  • 12:13 start acquisition on target (B-VRI sequence 3)

12:20 slew to STD SA95-275 (DDT Epifani) * 11:22 collimation and copointing * 12:28 start acquisition on SA95-275

12:47 sky flats LBC-R V_Bessel (PA=0)
Topic revision: r1 - 31 Jul 2024, UnknownUser
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback